# Graph Coloring for Multi-Task Learning

### v2

@@ D.5.2 Deterministic covariance sketching via frequent directions
We maintain a deterministic sketch $B\in\mathbb{R}^{\ell\times d}$ of the row space of $M$ using Frequent Directions and either project rows onto $\mathrm{span}(B)$ or form an approximate Gram from the sketch (Liberty, 2013; Ghashami et al., 2016). Maintaining the sketch costs $O(Kd\ell)$, the cosine Gram in the sketch space costs $O(K^{2}\ell)$, and storage for the sketch is $O(\ell d)$. Frequent Directions gives a spectral-norm bound
@@ D.5.3 Edge sampling for conflict graphs with adaptive refinement
We reduce the number of cosine evaluations by computing $\widehat{C}_{ij}$ for only $\tilde{O}(K\log K)$ randomly chosen task pairs to build a provisional conflict graph and then refining by evaluating additional pairs that are near the threshold or needed to certify connectivity and chromatic structure. We still compute all $K$ row norms once in $O(Kd)$ time for normalization, and the first pass costs $O(Kd\log K)$ for the sampled dot products. The total cost adds only the refinement work, which remains small when only few pairs are ambiguous. Under a planted separation model with margin $\gamma$ and reasonably dense cross-group conflicts, one can show with high probability that the sampled graph already captures the correct inter-group connectivity, so the coloring or component structure is recovered after the first pass and only boundary pairs need refinement. This reduces the pairwise work from $K^{2}$ to near $K\log K$ while preserving the final decisions under stated assumptions (Erdős & Rényi, 1960).
@@ Appendix F Convergence rate with $\tau$ -dependent constant
Because $F$ is LL–smooth, for any $\eta\leq\tfrac{1}{L}$ the standard non-convex SGD inequality (Ghadimi & Lan 2013, Lemma 3.2) gives
@@ Appendix G Bounded Staleness via Greedy Graph Coloring
iterations. In particular, the schedule enforces a bounded inter-update delay of at most $\Delta$ iterations per task, consistent with the bounded-delay assumption of Recht et al. (Niu et al., 2011).
@@
A greedy algorithm scans vertices in some order and assigns to each vertex the smallest available color not used by its already colored neighbors. When the $i$-th vertex $v$ is reached, at most $\deg(v)\leq\Delta$ of its neighbors are already colored, so at most $\Delta$ colors are unavailable. Therefore one of the first $\Delta+1$ colors is always free, implying $m\leq\Delta+1$ (Lovász, 2006).
@@ G.1 Interpretation
The bound (Equation 100) guarantees that the shared parameters used by any task are refreshed at least once every $\Delta$ iterations in the worst case (e.g., when the conflict graph is a clique of size $\Delta+1$). This aligns with the bounded-delay assumption common in analyses of asynchronous SGD and lock-free training, so convergence proofs built under that assumption apply to our cyclic schedule with delay parameter at most $\Delta$ when iterations are used as the unit of delay (Niu et al., 2011; Lian et al., 2015). In practice $\Delta$ is often much smaller than the total number of tasks, so the scheduler achieves low interference and low parameter staleness simultaneously.
@@ I.1 Proof of Theorem 8
The proof is a streamlined restatement of ((Ghadimi & Lan, 2013; Nemirovski et al., 2009)). By $L$-smoothness,
@@ I.2.1 Consistency with the unbiased SGD assumption
The NYU Depth Dataset v2 (NYUv2) (Silberman et al., 2012) consists of RGB-D indoor scenes with 1,449 densely labeled pairs of RGB and depth images. To demonstrate auxiliary task value in data-scarce conditions, we employ a subset of 250 training samples randomly selected from the original training set.
@@
All tasks utilize RGB images as the sole input modality, with depth maps and surface normals serving as prediction targets rather than input features. A ResNet-18 (He et al., 2015) backbone trained from scratch processes the RGB input, with task-specific decoder heads for segmentation (with 32 $\times$ upsampling), depth regression, surface normal regression, and color temperature estimation.
@@
The CIFAR-10 (Krizhevsky et al., 2009) dataset contains 60,000 $32\times 32$ color images across 10 generic classes. To evaluate our interference-aware scheduler in a data-scarce environment where auxiliary tasks provide maximum benefit, we employ a subset of 2,500 training samples (250 per class) from the original 50,000 training images.
For the multi-task learning setup, we set image classification as the main task and construct three auxiliary tasks synthetically from the RGB images. The positive auxiliary tasks include: (1) quadrant localization, where the model predicts which quadrant contains the primary object, and (2) texture classification using Gabor filter responses clustered into 8 texture categories via k-means clustering. The negative auxiliary tasks consist of: (3) corruption-type prediction, where images are artificially corrupted using 15 different corruption types from the ImageNet-C corruption suite (Hendrycks & Dietterich, 2019), and (4) rotation angle prediction, where images are rotated by 0°, 90°, 180°, or 270° and the model predicts the rotation angle.
All tasks share a ResNet-18 (He et al., 2015) backbone trained from scratch without pretraining, with task-specific heads for each auxiliary task.
@@
The AV‑MNIST benchmark (Vielzeuf et al., 2018) pairs MNIST images (Lecun et al., 1998) with a log‑mel spectrogram of the corresponding spoken digit from TIDIGITS (Leonard & Doddington, 1993). It is a synthetic benchmark that has significant noise applied to audio and feature reduction applied to images, making it far more difficult than the original MNIST.
We use all paired samples in our experiments. Our primary task is 10‑way digit classification. Following (Vielzeuf et al., 2018), we encode images with a small 4‑layer convolutional network and spectrograms with a 2‑layer CNN, both built and trained from scratch. These embeddings are projected and fused for processing by a simple MLP in intermediate fusion (Boulahia et al., 2021; Guarrasi et al., 2025), as are the models trained on MM-IMDb and STOCKS. We include only one positive auxiliary class, Digital Parity. This task aims to identify the digits as either even or odd, which has been shown to be a positive auxiliary task for improving representations on MNIST-like datasets (Tacchetti et al., 2018; Mohammadi et al., 2020).
@@
The MM‑IMDb dataset (Arevalo et al., 2017) contains 25,959 movies with genre annotations over 23 categories. We extract poster images and plot summaries for every movie in the dataset.
The images and summaries are encoded by a frozen VGG16 (Simonyan & Zisserman, 2014) and Google word2vec (Mikolov et al., 2013) model, respectively. Our main task is movie genre prediction. We add one positive auxiliary task, Release Decade, and one negative auxiliary task, the classification of the title’s first word as either a vowel or consonant.
@@
The STOCKS datasets we use, introduced in (Liang et al., 2021), contain stock market timeseries data across two categories. Specifically: (1) STOCKS-F&B, which has 14 input and 4 output stocks in the GICS Restaurants or Packaged Food & Meats category (MSCI Inc. & S&P Dow Jones Indices, 2024), and (2) STOCKS-HEALTH, which contains 56 input and 7 output stocks in the Health Care category.
Every input stock consists of 500 trading days, with the goal of predicting returns over the next day. We discretize the continuous return variable $R$ into three non-overlapping categories: (1) $Low$, where $0\leq R<0.1$, (2) $Medium$, where $0.1\leq R<0.5$, and (3) $High$, where $R\geq 0.5$. Mean Absolute Error (MAE) is calculated by mapping the three classes to numbers ($Low\rightarrow 0$, $Medium\rightarrow 1$, $High\rightarrow 2$) and then deriving MAE as usual. Each input series is encoded by the same CNN-BiLSTM network. This consists of 3 CNNs and 1 BiLSTM (Cui et al., 2018).
@@
    Gradnorm (Chen et al., 2018). Balances task learning rates by normalizing gradient magnitudes relative to target loss ratios. This maintains consistent training dynamics across tasks.
@@
    MGDA (Sener & Koltun, 2018). Formulates multi-task learning as a multi-objective optimization problem, finding Pareto-optimal solutions (Lockwood, 2008; Pareto, 2014) through gradient descent in the convex hull of gradients (Fliege & Svaiter, 2000; Miettinen, 1999).
@@
    PCGrad (Yu et al., 2020). Projects conflicting gradients onto orthogonal subspaces when negative cosine similarity is detected, eliminating destructive interference between task gradients.
@@
    CAGrad (Liu et al., 2021). Extends PCGrad by adaptively adjusting gradient magnitudes based on conflict severity. This proves more nuanced modifications to gradients than binary projection.
@@
    Adatask (Yang et al., 2023). Dynamically reweighs task losses using relative loss changes, adapting to varying task learning rates during training.
@@
    FAMO (Liu et al., 2023). Fast Adaptive Multitask Optimization dynamically adjusts task weights to equalize each task’s rate of loss improvement. It uses an online, per-step rule (no pairwise gradient ops), adding negligible overhead while remaining robust to loss-scale differences.
@@
    Fair Resource Allocation in MTL (FairGrad) (Ban & Ji, 2024). Views the shared update as a limited resource and chooses it to maximize an $\alpha$-fair utility of per-task improvements. The parameter $\alpha$ controls the trade-off between average performance and fairness.
@@
    Nash-MTL (Navon et al., 2022). Frames multitask training as a bargaining game and computes a scale-invariant weighted combination of task gradients given by the Nash bargaining solution. Weights are obtained by solving a small inner problem (e.g., via CCP) using the gradient Gram matrix. Updates are balanced across tasks.
@@
Multi-task learning (MTL) methods have evolved from simple loss-weighting approaches to larger and more sophisticated optimization techniques that manage task conflict and cooperation (Yang et al., 2023). Early adaptive-weighting approaches sought to balance losses automatically (Vandenhende et al., 2022; Fan et al., 2023), while more recent work modifies gradients directly (Yu et al., 2020). Task scheduling and grouping methods, though far less popular than adaptive weighting techniques (Torbarina et al., 2023), have contributed to the field by controlling the timing of updates.
@@
From early MTL work it became clear that simply summing task losses often favors one objective at the expense of others (Kurin et al., 2022; Zhao et al., 2024; Mueller et al., 2022), especially when losses have different scales or noise levels. To address this, practitioners manually tuned per‑task weight coefficients ($\lambda$-values) to rebalance learning (Argyriou et al., 2007; Ando & Zhang, 2005; Evgeniou et al., 2005; Kang et al., 2011; Liang & Zhang, 2020; Lin et al., 2022; Yu et al., 2021), but this process was laborious and dataset-specific. Thus, researchers began to develop automated methods.
@@
(Kendall et al., 2018) introduced uncertainty weighting, learning each task’s homoscedastic (constant‑variance) (Bishop, 2006) noise to scale losses automatically and improve depth and semantics on NYUv2 (Silberman et al., 2012).
GradNorm automatically balances multiple loss functions by tuning each task’s gradient magnitude so that all tasks train at comparable speeds (Chen et al., 2018). It does this by introducing a single asymmetry hyperparameter $\alpha$ that governs how much each task’s loss is scaled. This eliminates the need for expensive grid searches over manual weights. GradNorm was also a major leap empirically as it surpassed exhaustive search baselines on both regression and classification tasks. Dynamic Weight Averaging (DWA) extended this idea by adjusting weights based on loss rate of change, reducing oscillations between tasks (Liu et al., 2019).
More recently AdaTask applies task-specific learning rates that adapt to each head’s gradient norm, yielding significant gains on multi-label classification benchmarks (Yang et al., 2023).
@@
Rather than rescaling losses, gradient surgery methods alter update directions. PCGrad projects gradients that conflict (negative cosine) onto each other’s normal plane, significantly boosting efficiency on supervised vision and RL problems (Yu et al., 2020). CAGrad frames task balance as a min-max optimization, finding updates that maximize the worst-case task improvement (Liu et al., 2021). The Multiple Gradient Descent Algorithm (MGDA) computes a Pareto-optimal convex combination of task gradients, ensuring no task is harmed (Sener & Koltun, 2018). More recent variants such as SAM-GS incorporate momentum into conflict detection, smoothing gradient estimates while preserving the benefits of surgery (Borsani et al., 2025).
M.4 Task Grouping
@@
Early approaches under this category used round-robin and random sampling-based approaches that ignored any task relationships (McCann et al., 2018; Zamir et al., 2020). Standley et al. (2020) exhaustively searches over small subsets to identify beneficial groupings, demonstrating the potential of selective updates but failing to scale beyond eight tasks due to computational complexity.
@@
Task Affinity Groupings (TAG) (Fifty et al., 2021) performs one joint training run to measure inter-task ’affinity’. It quantifies how an update for task $i$ (its gradient) would change task $j$’s loss, and it uses these cross-effects to select partitions of tasks that should share updates. The key idea is to treat grouping as an outcome of measured gradient interactions.
@@
Ayman et al. (Ayman et al., 2023) train a predictor that maps single-task statistics and dataset features to an estimate of whether two or more tasks should be grouped. They then use that predictor to guide a randomized search over groups, which dramatically reduces the number of multi-task trainings (or ’MTL trials’) needed to find a good partition.
@@
Using a completely different approach, Towards Principled Task Grouping (PTG) (Wang et al., 2024) formulates grouping as a mathematical program with a theoretically motivated objective capturing beneficial transfer while respecting resource constraints (e.g., compute budgets). It builds a principled optimization over candidate groups that is meant to generalize across application domains.

### v4

@@ D.5.2 Deterministic covariance sketching via frequent directions
We maintain a deterministic sketch $B\in\mathbb{R}^{\ell\times d}$ of the row space of $M$ using Frequent Directions and either project rows onto $\mathrm{span}(B)$ or form an approximate Gram from the sketch (Liberty2013; Ghashami2016). Maintaining the sketch costs $O(Kd\ell)$, the cosine Gram in the sketch space costs $O(K^{2}\ell)$, and storage for the sketch is $O(\ell d)$. Frequent Directions gives a spectral-norm bound
@@ D.5.3 Edge sampling for conflict graphs with adaptive refinement
We reduce the number of cosine evaluations by computing $\widehat{C}_{ij}$ for only $\tilde{O}(K\log K)$ randomly chosen task pairs to build a provisional conflict graph and then refining by evaluating additional pairs that are near the threshold or needed to certify connectivity and chromatic structure. We still compute all $K$ row norms once in $O(Kd)$ time for normalization, and the first pass costs $O(Kd\log K)$ for the sampled dot products. The total cost adds only the refinement work, which remains small when only few pairs are ambiguous. Under a planted separation model with margin $\gamma$ and reasonably dense cross-group conflicts, one can show with high probability that the sampled graph already captures the correct inter-group connectivity, so the coloring or component structure is recovered after the first pass and only boundary pairs need refinement. This reduces the pairwise work from $K^{2}$ to near $K\log K$ while preserving the final decisions under stated assumptions (ErdosRenyi1960).
@@ D.5.4 Incremental gram updates
D.5.5 Experimental Analysis
Table 3: Runtime and Taskonomy Tiny validation metrics for SON-GOKU scheduling variants.
Family Variant Runtime Taskonomy Tiny
Elapsed (s) $\downarrow$ Imgs/s $\uparrow$ Refresh (ms) $\downarrow$ Depth Eucl. RMSE $\downarrow$ Depth Eucl. MAE $\downarrow$ Normal mean (deg) $\downarrow$ Reshading MAE $\downarrow$
[…]
We evaluated each optimizations’ speed and relative impact on performance on the Taskonomy Tiny subset. We use the Tiny subset here as this experiment does not require large-scale training for valid results.
Results are presented in Table 3. We can see that, in practice, every approach actually does improve speed over the SON-GOKU baseline (Frequent Directions with high $r$, note that we use $r$ in the same way that ghashami2016frequent use $\ell$). However, every approach also degrades performance (by performance we are referring the main objective metric, like loss or accuracy, not speed) to varying extents [^1] Please note that varying performance for the SON-GOKU baseline across experiments is due to changes in the backbone model architecture and the Taskonomy subset.. Edge sampling and incremental gram updates have an extremely negative impact on performance. Such approaches may need additional fine-tuning or may only be practical in specific settings. Interestingly, we observe a slight decrease in performance when using lower width for Frequent Directions, demonstrating the tradeoffs between speed and performance that come with such an approach.
Overall, every approach achieves the desired effect of increasing speed, but does so at varying costs in performance. Each approach will require careful fine tuning and usage in real-world deployment to properly weigh this tradeoff. Based on our experimentation, it appears that Frequent Directions offers the most consistent and reasonable tradeoff that clearly scales with the FD width.
@@ Appendix F Convergence rate with $\tau$ -dependent constant
Because $F$ is LL–smooth, for any $\eta\leq\tfrac{1}{L}$ the standard non-convex SGD inequality (GhadimiLan2013, Lemma 3.2) gives
@@ Appendix G Bounded Staleness via Greedy Graph Coloring
iterations. In particular, the schedule enforces a bounded inter-update delay of at most $\Delta$ iterations per task, consistent with the bounded-delay assumption of Recht et al. (Recht2011Hogwild).
@@
A greedy algorithm scans vertices in some order and assigns to each vertex the smallest available color not used by its already colored neighbors. When the $i$-th vertex $v$ is reached, at most $\deg(v)\leq\Delta$ of its neighbors are already colored, so at most $\Delta$ colors are unavailable. Therefore one of the first $\Delta+1$ colors is always free, implying $m\leq\Delta+1$ (lovasz2006graph).
@@ G.1 Interpretation
The bound (Equation 101) guarantees that the shared parameters used by any task are refreshed at least once every $\Delta$ iterations in the worst case (e.g., when the conflict graph is a clique of size $\Delta+1$). This aligns with the bounded-delay assumption common in analyses of asynchronous SGD and lock-free training, so convergence proofs built under that assumption apply to our cyclic schedule with delay parameter at most $\Delta$ when iterations are used as the unit of delay (Recht2011Hogwild; lian2015asynchronous). In practice $\Delta$ is often much smaller than the total number of tasks, so the scheduler achieves low interference and low parameter staleness simultaneously.
@@ I.1 Proof of Theorem 8
The proof is a streamlined restatement of ((GhadimiLan2013; Nemirovski2009)). By $L$-smoothness,
@@ I.2.1 Consistency with the unbiased SGD assumption
The NYU Depth Dataset v2 (NYUv2) (Silberman2012) consists of RGB-D indoor scenes with 1,449 densely labeled pairs of RGB and depth images. To demonstrate auxiliary task value in data-scarce conditions, we employ a subset of 250 training samples randomly selected from the original training set.
@@
All tasks utilize RGB images as the sole input modality, with depth maps and surface normals serving as prediction targets rather than input features. A ResNet-18 (he2016residual) backbone trained from scratch processes the RGB input, with task-specific decoder heads for segmentation (with 32 $\times$ upsampling), depth regression, surface normal regression, and color temperature estimation.
@@
The CIFAR-10 (krizhevsky2009learning) dataset contains 60,000 $32\times 32$ color images across 10 generic classes. To evaluate our interference-aware scheduler in a data-scarce environment where auxiliary tasks provide maximum benefit, we employ a subset of 2,500 training samples (250 per class) from the original 50,000 training images.
For the multi-task learning setup, we set image classification as the main task and construct three auxiliary tasks synthetically from the RGB images. The positive auxiliary tasks include: (1) quadrant localization, where the model predicts which quadrant contains the primary object, and (2) texture classification using Gabor filter responses clustered into 8 texture categories via k-means clustering. The negative auxiliary tasks consist of: (3) corruption-type prediction, where images are artificially corrupted using 15 different corruption types from the ImageNet-C corruption suite (hendrycks2019robustness), and (4) rotation angle prediction, where images are rotated by 0°, 90°, 180°, or 270° and the model predicts the rotation angle.
All tasks share a ResNet-18 (he2016residual) backbone trained from scratch without pretraining, with task-specific heads for each auxiliary task.
@@
The AV‑MNIST benchmark (vielzeuf2019centralnet) pairs MNIST images (lecun1998gradient) with a log‑mel spectrogram of the corresponding spoken digit from TIDIGITS (leonard1993tidigits). It is a synthetic benchmark that has significant noise applied to audio and feature reduction applied to images, making it far more difficult than the original MNIST.
We use all paired samples in our experiments. Our primary task is 10‑way digit classification. Following (Vielzeuf et al., 2018), we encode images with a small 4‑layer convolutional network and spectrograms with a 2‑layer CNN, both built and trained from scratch. These embeddings are projected and fused for processing by a simple MLP in intermediate fusion (boulahia2021early; guarrasi2025systematic), as are the models trained on MM-IMDb and STOCKS. We include only one positive auxiliary class, Digital Parity. This task aims to identify the digits as either even or odd, which has been shown to be a positive auxiliary task for improving representations on MNIST-like datasets (Tacchetti2018; Mohammadi2020OddOneOut).
@@
The MM‑IMDb dataset (arevalo2017gatedmultimodalunitsinformation) contains 25,959 movies with genre annotations over 23 categories. We extract poster images and plot summaries for every movie in the dataset.
The images and summaries are encoded by a frozen VGG16 (simonyan2014very) and Google word2vec (mikolov2013efficientestimationwordrepresentations) model, respectively. Our main task is movie genre prediction. We add one positive auxiliary task, Release Decade, and one negative auxiliary task, the classification of the title’s first word as either a vowel or consonant.
@@
The STOCKS datasets we use, introduced in (liang2021multibench), contain stock market timeseries data across two categories. Specifically: (1) STOCKS-F&B, which has 14 input and 4 output stocks in the GICS Restaurants or Packaged Food & Meats category (MSCI2024GICs), and (2) STOCKS-HEALTH, which contains 56 input and 7 output stocks in the Health Care category.
Every input stock consists of 500 trading days, with the goal of predicting returns over the next day. We discretize the continuous return variable $R$ into three non-overlapping categories: (1) $Low$, where $0\leq R<0.1$, (2) $Medium$, where $0.1\leq R<0.5$, and (3) $High$, where $R\geq 0.5$. Mean Absolute Error (MAE) is calculated by mapping the three classes to numbers ($Low\rightarrow 0$, $Medium\rightarrow 1$, $High\rightarrow 2$) and then deriving MAE as usual. Each input series is encoded by the same CNN-BiLSTM network. This consists of 3 CNNs and 1 BiLSTM (Cui2018).
@@
    Gradnorm (Chen2017GradNorm). Balances task learning rates by normalizing gradient magnitudes relative to target loss ratios. This maintains consistent training dynamics across tasks.
@@
    MGDA (Sener2018). Formulates multi-task learning as a multi-objective optimization problem, finding Pareto-optimal solutions (lockwood2008pareto; pareto2014manual) through gradient descent in the convex hull of gradients (FliegeSvaiter2000; Miettinen1999).
@@
    PCGrad (Yu2020PCGrad). Projects conflicting gradients onto orthogonal subspaces when negative cosine similarity is detected, eliminating destructive interference between task gradients.
@@
    CAGrad (Liu2021CAGrad). Extends PCGrad by adaptively adjusting gradient magnitudes based on conflict severity. This proves more nuanced modifications to gradients than binary projection.
@@
    Adatask (Yang2023AdaTask). Dynamically reweighs task losses using relative loss changes, adapting to varying task learning rates during training.
@@
    FAMO (Liu2023FAMO). Fast Adaptive Multitask Optimization dynamically adjusts task weights to equalize each task’s rate of loss improvement. It uses an online, per-step rule (no pairwise gradient ops), adding negligible overhead while remaining robust to loss-scale differences.
@@
    Fair Resource Allocation in MTL (FairGrad) (Ban2024FairMTL). Views the shared update as a limited resource and chooses it to maximize an $\alpha$-fair utility of per-task improvements. The parameter $\alpha$ controls the trade-off between average performance and fairness.
@@
    Nash-MTL (Navon2022NashMTL). Frames multitask training as a bargaining game and computes a scale-invariant weighted combination of task gradients given by the Nash bargaining solution. Weights are obtained by solving a small inner problem (e.g., via CCP) using the gradient Gram matrix. Updates are balanced across tasks.
K.3 Ablation Study Models
  1. 1.
    Static One-Shot Coloring. We run the greedy graph coloring once at the start of training, freeze the resulting task groups, and never recompute the conflict graph. All other hyperparameters ($\tau$, history length $H$, and update interval $R$) match the full scheduler. As training progresses we expect the fixed coloring to grow stale, mixing tasks whose interference relationships have changed. This ablation isolates the benefit of dynamic recoloring, showing how much performance depends on adapting the schedule to evolving gradient conflicts.
  2. 2.
    Single-Step Conflict Estimation. Here, we set the history length to $H=1$, so every recoloring step relies on only the most recent mini-batch gradients to estimate interference. Without aggregation over many past steps, the conflict graph should become highly noisy, causing unstable task groupings from one update window to the next. This variant tests the importance of historical conflict statistics in the scheduler. Threshold Graph (baseline). We connect tasks $i$ and $j$ whenever the smoothed cosine $\hat{s}_{ij}(t)$ falls below a global threshold $\tau(t)$. This is the rule used in the main method and analyzed in our recovery and scheduling theory. It prioritizes the most strongly conflicting pairs globally and serves as the reference against which the other graph rules are compared.
  3. 3.
    kNN-Symmetric Graph. For each task we identify its $m$ most conflicting neighbors (those with the smallest smoothed cosine values) and add edges to those neighbors. We then symmetrize the graph by including an undirected edge if either endpoint selects the other. This construction roughly fixes the degree of each node and spreads edges more evenly across tasks. It tests whether enforcing local degree control can outperform or match the global threshold rule when conflict is heterogeneous across tasks.
  4. 4.
    Signed-Only Graph. Here we connect two tasks only if their smoothed cosine is strictly negative, $\hat{s}_{ij}(t)<0$, ignoring the magnitude of the conflict. This yields a much sparser graph that only records clearly antagonistic pairs. This ablation explores an extreme notion of conflict and allows us to test whether discarding moderately conflicting (but still harmful) interactions degrades performance.
  5. 5.
    Quantile Threshold Graph. Instead of fixing $\tau$ by hand, we set $\tau(t)$ at each refresh so that only the worst $p\%$ of smoothed cosine values are treated as conflicting. This behaves like an adaptive threshold that tracks how the similarity distribution drifts over training, keeping the graph density approximately stable over time. This variant tests whether such an automatically tuned cutoff provides an advantage over the fixed global threshold.
We evaluate each graph rule under two settings. In the fixed $\tau$ setting, all rules share the same $\tau(t)$ schedule used in the main experiments, and we simply observe the induced edge density, number of colors, and validation performance. In the density-matched setting, we adjust the hyperparameters of each rule (e.g., $m$ for kNN, percentile $p$ for quantile) so that all graphs have approximately the same edge density at each refresh. This isolates the effect of which pairs are marked as conflicting, rather than how many edges are present.
@@
Multi-task learning (MTL) methods have evolved from simple loss-weighting approaches to larger and more sophisticated optimization techniques that manage task conflict and cooperation (Yang2023AdaTask). Early adaptive-weighting approaches sought to balance losses automatically (9336293; Fan2023MaxGNR), while more recent work modifies gradients directly (Yu2020PCGrad). Task scheduling and grouping methods, though far less popular than adaptive weighting techniques (torbarina2023challengesopportunitiesusingtransformerbased), have contributed to the field by controlling the timing of updates.
@@
From early MTL work it became clear that simply summing task losses often favors one objective at the expense of others (Kurin2022InDefense; Zhao2024ExcessMTL; Mueller2022Temperature), especially when losses have different scales or noise levels. To address this, practitioners manually tuned per‑task weight coefficients ($\lambda$-values) to rebalance learning (Argyriou2007MultiTaskFeature; AndoZhang2005; Evgeniou2005; Kang2011; Liang2020; lin2022a; Yu2021RevisitingLossWeight), but this process was laborious and dataset-specific. Thus, researchers began to develop automated methods.
@@
(kendall2018multi) introduced uncertainty weighting, learning each task’s homoscedastic (constant‑variance) (Bishop2006) noise to scale losses automatically and improve depth and semantics on NYUv2 (Silberman2012).
GradNorm automatically balances multiple loss functions by tuning each task’s gradient magnitude so that all tasks train at comparable speeds (Chen2017GradNorm). It does this by introducing a single asymmetry hyperparameter $\alpha$ that governs how much each task’s loss is scaled. This eliminates the need for expensive grid searches over manual weights. GradNorm was also a major leap empirically as it surpassed exhaustive search baselines on both regression and classification tasks. Dynamic Weight Averaging (DWA) extended this idea by adjusting weights based on loss rate of change, reducing oscillations between tasks (Liu2019EndToEnd).
More recently AdaTask applies task-specific learning rates that adapt to each head’s gradient norm, yielding significant gains on multi-label classification benchmarks (Yang2023AdaTask).
@@
Rather than rescaling losses, gradient surgery methods alter update directions. PCGrad projects gradients that conflict (negative cosine) onto each other’s normal plane, significantly boosting efficiency on supervised vision and RL problems (Yu2020PCGrad). CAGrad frames task balance as a min-max optimization, finding updates that maximize the worst-case task improvement (Liu2021CAGrad). The Multiple Gradient Descent Algorithm (MGDA) computes a Pareto-optimal convex combination of task gradients, ensuring no task is harmed (Sener2018). More recent variants such as SAM-GS incorporate momentum into conflict detection, smoothing gradient estimates while preserving the benefits of surgery (Borsani2025SAMGS).
M.4 Empirical Task Grouping
@@
Early approaches under this category used round-robin and random sampling-based approaches that ignored any task relationships (McCann2018NaturalLanguageDecathlon; Zamir2020RobustLearning). Standley2020TaskGrouping exhaustively searches over small subsets to identify beneficial groupings, demonstrating the potential of selective updates but failing to scale beyond eight tasks due to computational complexity.
@@
Task Affinity Groupings (TAG) (Fifty2021) performs one joint training run to measure inter-task ’affinity’. It quantifies how an update for task $i$ (its gradient) would change task $j$’s loss, and it uses these cross-effects to select partitions of tasks that should share updates. The key idea is to treat grouping as an outcome of measured gradient interactions.
@@
Ayman et al. (ayman2023taskgroupingautomatedmultitask) train a predictor that maps single-task statistics and dataset features to an estimate of whether two or more tasks should be grouped. They then use that predictor to guide a randomized search over groups, which dramatically reduces the number of multi-task trainings (or ’MTL trials’) needed to find a good partition.
@@
Using a completely different approach, Towards Principled Task Grouping (PTG) (wang2024towards) formulates grouping as a mathematical program with a theoretically motivated objective capturing beneficial transfer while respecting resource constraints (e.g., compute budgets). It builds a principled optimization over candidate groups that is meant to generalize across application domains.